# Investigating the Impact of Model Instability on Explanations and Uncertainty

### v1

@@ D.4 Uncertainty and explanation plausibility at high levels of perturbation
Predictive Uncertainty Epistemic Uncertainty Dataset Model GBP IXG IG SG GBP IXG IG SG SST-2 BERT -0.016 0.020 -0.015 0.092 -0.162 -0.100 -0.089 -0.011 ELECTRA -0.122 -0.114 -0.048 -0.032 -0.308 -0.289 -0.160 -0.151 RoBERTa -0.169 -0.123 -0.153 -0.130 -0.315 -0.254 -0.244 -0.178 GPT2 -0.075 -0.017 -0.070 -0.016 -0.159 -0.100 -0.096 -0.048 OPT -0.013 -0.013 -0.054 -0.054 -0.070 -0.070 -0.020 -0.020 SemEval BERT 0.088 0.103 0.088 0.103 0.089 0.104 0.087 0.103 ELECTRA 0.103 0.096 0.103 0.096 0.105 0.097 0.104 0.097 RoBERTa 0.106 0.106 0.106 0.106 0.108 0.106 0.104 0.104 GPT2 0.064 0.083 0.065 0.083 0.065 0.085 0.065 0.084 OPT 0.074 0.074 0.073 0.073 0.067 0.067 0.076 0.076 HateXplain BERT -0.049 -0.078 -0.049 -0.078 -0.040 -0.060 -0.041 -0.064 ELECTRA -0.054 -0.084 -0.061 -0.091 -0.033 -0.059 -0.060 -0.090 RoBERTa -0.021 -0.054 -0.023 -0.055 -0.009 -0.036 -0.020 -0.052 GPT2 0.134 0.090 0.140 0.094 0.126 0.097 0.134 0.092 OPT 0.019 0.019 0.003 0.003 -0.025 -0.025 0.005 0.005

### v2

@@ D.4 Uncertainty and explanation plausibility at high levels of perturbation
{tblr} roweven = r, row3 = r, row5 = r, row7 = r, row9 = r, row11 = r, row13 = r, row15 = r, row17 = r, cell13 = c=4c, cell17 = c=4c, cell31 = r=5, cell81 = r=5, cell131 = r=5, vline4,8 = 1, vline3,7,11 = 2-17, hline1,18 = -0.08em, hline2-3,8,13 = -0.05em, & Predictive Uncertainty Epistemic Uncertainty
Dataset Model GBP IXG IG SG GBP IXG IG SG
SST-2 BERT -0.016 0.020 -0.015 0.092 -0.162 -0.100 -0.089 -0.011
ELECTRA -0.122 -0.114 -0.048 -0.032 -0.308 -0.289 -0.160 -0.151
RoBERTa -0.169 -0.123 -0.153 -0.130 -0.315 -0.254 -0.244 -0.178
GPT2 -0.075 -0.017 -0.070 -0.016 -0.159 -0.100 -0.096 -0.048
OPT -0.013 -0.013 -0.054 -0.054 -0.070 -0.070 -0.020 -0.020
SemEval BERT 0.088 0.103 0.088 0.103 0.089 0.104 0.087 0.103
ELECTRA 0.103 0.096 0.103 0.096 0.105 0.097 0.104 0.097
RoBERTa 0.106 0.106 0.106 0.106 0.108 0.106 0.104 0.104
GPT2 0.064 0.083 0.065 0.083 0.065 0.085 0.065 0.084
OPT 0.074 0.074 0.073 0.073 0.067 0.067 0.076 0.076
HateXplain BERT -0.049 -0.078 -0.049 -0.078 -0.040 -0.060 -0.041 -0.064
ELECTRA -0.054 -0.084 -0.061 -0.091 -0.033 -0.059 -0.060 -0.090
RoBERTa -0.021 -0.054 -0.023 -0.055 -0.009 -0.036 -0.020 -0.052
GPT2 0.134 0.090 0.140 0.094 0.126 0.097 0.134 0.092
OPT 0.019 0.019 0.003 0.003 -0.025 -0.025 0.005 0.005